OpenStudy (anonymous):

If (a,-8) is a point on the graph of y=x^2-6 what is a?

jimthompson5910 (jim_thompson5910):

y=x^2-6 -8=a^2-6 ... plug in x = a and y = -8 now solve for 'a'